40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 10 15:41:37 2019, aaron, Update, IOO, WFS measurements Screenshot_from_2019-09-10_18-51-28.pngmctrend_190910.png
    Reply  Wed Sep 11 10:26:56 2019, aaron, Update, IOO, WFS measurements 
       Reply  Wed Sep 11 14:37:43 2019, aaron, Update, IOO, WFS measurements 20190911_WFS.jpg20190911_WFS_2.jpg
          Reply  Thu Sep 12 12:42:31 2019, aaron, Update, IOO, WFS measurements WFS2_TFs.pdfD755499D-9FDF-4E2B-BFC1-016B459DD35D.jpeg
             Reply  Fri Sep 13 10:36:03 2019, aaron, Update, IOO, WFS measurements TF-.pngWFS2_TFs.pdf
                Reply  Mon Sep 16 12:38:59 2019, aaron, Update, IOO, WFS measurements 
                Reply  Thu Oct 3 11:38:35 2019, aaron, Update, IOO, WFS measurements 
          Reply  Mon Sep 30 13:20:55 2019, aaron, Update, IOO, shot noise measurement 
       Reply  Fri Sep 13 10:53:40 2019, aaron, Update, IOO, WFS loop measurements 
          Reply  Mon Sep 16 05:08:04 2019, rana, Update, IOO, WFS loop measurements LSCsensingMatrix.ipynbASCsensingMatrix.ipynb
             Reply  Mon Sep 16 11:55:58 2019, rika, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 09:41:48 2019, gautam, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 10:47:44 2019, rika, Update, IOO, WFS loop measurements MC2.pdf
                   Reply  Wed Sep 18 14:45:52 2019, rika, Update, IOO, WFS loop measurements Screenshot_from_2019-09-18_18-15-34.png
             Reply  Tue Sep 17 10:34:48 2019, aaron, Update, IOO, WFS loop measurements 
                Reply  Tue Oct 8 20:39:42 2019, aaron, Update, IOO, WFS loop measurements 
                   Reply  Wed Oct 9 09:37:28 2019, aaron, Update, IOO, WFS loop measurements 
Message ID: 14929     Entry time: Thu Oct 3 11:38:35 2019     In reply to: 14875
Author: aaron 
Type: Update 
Category: IOO 
Subject: WFS measurements 

I set up the spectrum analyzer to make the WFS head RF transfer function measurement (V/W) on WFS1. I placed the Jenne laser on the AP table, along with the reference PD power supply, laptop, and laser power supply. The Agilent output AM modulates the laser; the reference PD is again NewFocus 1611, with its AC output sent to Agilent's R channel and DC output sent to an oscilloscope;

At Koji's suggestion, I've started setting up a small breadboard to hold the fiber collimator, BS, and reference PD. I haven't really used fiber optics before, I'd appreciate another set of eyes before I get too deep.
Gautam showed me the collimator and fiber BS.

I closed the PSL shutter while checking for a location to place the breadboard, and opened it while writing this. Headed back to Cryo to pick up the large incandescent bulb we'd borrowed over the summer.

ELOG V3.1.3-